find the center and radius of the circle with the given equation:

x^2 + y^2 - 12x + 2y + 21 = 0

Please show detailed step-by-step directions so I can figure out how to do this on my own, Thank you!
Answer #1

Given,

Equation is x^{2}+y^{2}-12x+2y+21=0

General equation of circle is x^{2}+y^{2}+2gx+2fy+c=0

Center is C=(-g,-f)

Radius is r=\sqrt{g^{2}+f^{2}-c}

Now comparing x^{2}+y^{2}-12x+2y+21=0 with x^{2}+y^{2}+2gx+2fy+c=0

Here,

2g=-12

g=-6

2f=2

f=1

Now Center C will be, C=(-g,-f)=(6,-1)

Radius will be,

r=\sqrt{g^{2}+f^{2}-c}

r=\sqrt{(-6)^{2}+1^{2}-21}

r=\sqrt{36+1-21}

r=\sqrt{37-21}

r=\sqrt{16}

r=4

Therefore,

Center=(6,-1)

Radius = 4 units